News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 11 August 1999 7.00am EST

The following is an overview of the events that occurred over the last 24 hours

Completed Gyro Low Rate Bias Determination

Continued ACIS Background Measurements

Optical Cavity pressure measurement reading was 7.8E-06 compared with 8.1, which had been holding at 8.1 for the past 48 hours.

Complete Earth Scan #24.

Events scheduled for the next 24 hours
  • Open Aft Contamination Cover and begin taking data with ACIS from Flight Contamination Monitors (FCM) that are located on forward HRMA contamination door.
  • Begin ACA Dark Current Calibration.
  • Activate the HRC anti co-incidence shield.
  • Attempt to turn on HRC high voltage to full voltage






Return to Status Reports